OpenStudy (anonymous):

Weights of male mountain lions follow the normal distribution with a median of 140 lb and an interquartile range of 8 lb. How to find the 55th percentile??

OpenStudy (anonymous):

How to find the 75th and 95 thpercentile of the weights.

OpenStudy (anonymous):

@kniedite any idea?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

@kniedite since Iq=Up-Lq=8.2, median is given 150lb , how we can find lower quartile?

OpenStudy (anonymous):

a.154.11 b.160

OpenStudy (anonymous):

Q3-Q1= SD(z3-z1) 8.2 = SD(0.67-(-0.67)) SD=8.2/1.34 SD=6.12 0.674*6.12 + 150 =154.11 for 75th percentile 1.645*6.12 + 150 = 160 for 95th percentile